# code-for-first-row = , code-for-last-row = , code-for-first-col = , code-for-last-col =

### v1

@@ 3.3 Natural Language Processing
Compositionality[220] which describes the ability to easily recombine familiar units like words into novel phrases and sentences has been the focus of intense interest in artificial intelligence in recent years. However, existing neural network based models have proven to be extremely deficient in such a capability[308]. In order to improve on the compositional generalization ability of neural encoder-decoder architectures for natural language understanding, Guo et al. [105] introduce a hierarchical poset decoding paradigm with the goal of taking into consideration the partial permutation invariance and the poset structure of semantics. Compositional generalization is important in tasks that can involve complex natural language utterances, which can be associated to many equivalent meaning representations. One of such task is dependency parsing. Semantic dependency parsing[221] is the task of mapping natural language sentences into representations of their meaning in the form of directed graphs on words. Dependency parsing is generally used to identify semantic relations between words in a sentence by analyzing their grammatical structure. Dyer [83] proposes a novel method of for converting dependency trees to surface orders via syntactic word embeddings and edge-weighted posets. Graph neural network were employed to learn the edge weights of the poset representing the surface order and the training was carried out on the Universal Dependencies (UD) corpora. [222]
@@ 3.4 Classification
Most major studies in artificial intelligence has concentrated significantly on developing mechanisms for classification. Most prediction tasks in machine learning are considered as classification when the target variable is discrete. It is common to find learning tasks that involve ranking of objects particularly in domains that generate relational data where the goal of inferring is conditioned on a particular target object . Classification is the task of associating a single label $y\in Y$ with a covariate $x$. A generalization of this problem is conditional ranking, the task of assigning to $x$ a full or partial ranking of the items in Y. Lebabnon and Lafferty[129] propose a unifying algebraic framework for ensemble methods of classification and conditional ranking, focusing on the case of discrete and ranked inputs. Their approach is based on the ranking poset having a fixed size of items, which consists of the collection of all full and partial rankings equipped with the partial order given by refinement of rankings. The structure of the poset of partial ranking over Y gives rise to natural invariant distance functions that generalize Kendall’s Tau[309] and the Hamming distance. Using the invariance properties of the distances, a generative interpretation of the probabilistic model is derived, and this is suggested to be useful in model selection and validation[129]. Zhang [152] presents a framework for ranking with classification rules. Classification rules can be generated using inductive learning algorithms on data [226]. Sahami [152] introduces an inductive machine learning algorithm called Ruleamer which roduces a set of symbolic classification rules as output using lattices.The Ruleamer algorithm takes as input (i) a lattice, L, (ii) a set of instance classification labelings, C, which correspond to the instance nodes in L, and (iii) a noise parameter, N, indicating a percentage by which each induced rule can misclassify some portion of the training instances to which it applies. Learning classification rules has also been extended to sequence data [224, 225]. The problem of sequential classification is considered when classification states form lattices and response distributions are experiment-specific. [172] proposes a Bayesian framework for sequential classification on finite lattice models in which response distributions are allowed to vary according to experiment When the models are latent and complex, such as in cognitive applications, it becomes imperative to have available a variety of data analytic tools for fitting the models, and for the validation of assumptions that are made regarding the specification of class conditional response distributions for experiments. A framework for model fitting and analysis of experiments has been developed byTatsuoka[165]. This led to the development of a general theorem in [166, 167] on the asymptotically optimal sequential selection of experiments which is presented and applied to a Bayesian classification problem when the parameter space is a finite partially ordered. In general, lattices can be used as classification models when it is natural to consider classification states as partially ordered. Examples include cognitive modeling, such as in neuropsychological (NP) assessment, educational testing, and the group testing data.
3.5 Deep Unsupervised Learning
@@
Deep unsupervised learning[310] is an emerging discipline in the field of artificial intelligence and machine learning. It can serve as an alternative to supervised approaches in representation learning. In particular, unlabelled data is much cheaper to obtain which is a key advantage when adopting deep unsupervised learning for complex processing tasks involving large datasets. One of such task is visual similarities learning[227] in the field of computer vision. Various approaches have been adopted for visual similarity learning such as the use of convolution neural networks(CNN) [228]. But this approach has a major limitation in that it requires millions of samples of supervised training in order to achieve a good performance. Bautista et al.[20], propose an unsupervised approach to similarity learning based on CNNs by framing it as a combination of surrogatee (i.e. artificially created) classification tasks and poset ordering. This approach which combines classification with a partial ordering of samples solved some of the major problems of visual similarity learning such as:
@@ 3.6 Semi-Supervised Learning
Semi-supervised learning[229] is a branch of machine learning that bridges supervised and unsupervised learning by combining both labeled and unlabeled data to train artificial intelligence (AI) models for classification or regression. Semi-supervised classification methods are particularly relevant to scenarios where labelled data is scarce, expensive or difficult to obtain. Graph-based semi-supervised learning methods[230] have demonstrated their advantages in various domains due to their uniqueness of structure, the universality of applications, and their scalability to large scale data. Graph-based learning methods explicitly consider the relations between two entities (i.e.vertices). In an ordinary graph an edge connects exactly two vertices. On the other hand, a hypergraph represents a generalization of a graph in which an edge can join any number of vertices. But in most settings hypergraph structure represent the hyperedge as an unordered set of vertices without taking into consideration the possibility of ordering relationships among vertices which is inherent in most real-world data. This motivated Feng et al. [99] to suggest two related solutions which are:
  • the construction of a new data structure named Partial-Order Hypergraph, which specifically injects the partially ordering relations among vertices into a hyperedge ;
@@
In particular, graph gonvolution network(GCN)[312] was applied in a semi-supervied setting for learning over the partial-order hypergraphs in [99].
@@ 3.7 Time Series Modeling
Many machine learning and statistical models have been suggested for time series modeling of sequential data [313, 314]. One way of generating sequence data is from a series of discrete events that occur over time. Event sequence data arise in many applications, such as web browsing, e-commerce, and monitoring of processes. An important problem in mining sets of sequences of events is to get an overview of the ordering relationships in the data. [134] introduce a method for discovering partial orders from a sequence of events. The main idea of the method is to view a partial order as a generative model for sequences and by describing a set of sequences by means of mixture models of partial orders. In addition, it is also shown that the likelihood of a given partial order is inversely proportional to the number of total orders compatible with the partial order. The computation of the number of linear extensions was restricted to series paprallel posets which are computationally efficient.
Nicholls et al. [288] studies Bayesian inference for partial orders from random linear extensions using time-series data of social hierarchy during the time period duringthe 12th century. The outcome of the approach a new class of models for rank-order time series data in which actors are listed in order of precedence are presented as a collection of lists. The lists are modeled as a realisation of a queue in which queue-position is constrained by an underlying social hierarchy corresponding to a partial order in which the lists are random linear extensions. A consistent prior for the stochastic process of partial orders based on a latent variable representation for the partial order Model fitting to witness lists from Royal Acta from England, Wales and Normandy in the eleventh and twelfth centuries.
@@ 3.8 Learning to rank
Learning to rank (LTR)[315] describes a class of algorithmic techniques that apply supervised machine learning to solve ranking problems from a wide range of applications in domains such as information retrieval, recommendation systems, search engine optimization etc. It arises from the need to obtain information in the order of relevance during retrieval of a prediction task. Several methods have been proposed for LTR and they can be categorized or grouped into three main approaches: pointwise, pairwise , listwise. Pointwise approaches look at a single document at a time in the loss function. They essentially take a single document and train a classifier / regressor on it to predict how relevant it is for the current query. Pairwise approaches look at a pair of documents at a time in the loss function. Given a pair of documents, they try and come up with the optimal ordering for that pair and compare it to the ground truth. The goal for the ranker is to minimize the number of inversions in ranking i.e. cases where the pair of results are in the wrong order relative to the ground truth. Listwise approaches directly look at the entire list of documents and try to come up with the optimal ordering for it.
@@
In [268], it is noted that not all rankings can follow a strict total order in the context of learning to rank. As such a proposed method of relaxing the conventional setting such that predictions are given in terms of partial instead of total orders is outlined. The key idea is that if a model is uncertain about the relative order of two alternatives, in which case it is unable to clearly determine whether the former should precede the latter and vice-versar, it may decline from making a decison and instead declare such pair of alternatives as incomparable. Such learning to ranking problem can be formulated as follows: Given a set of training instances $\{x_{1},...,x_{n}\}\subseteq\mathcal{X}$ and a set of labels $\mathcal{Y}=\{y_{1},...,y_{k}\}$ endowed with an order $y_{1}<y_{2}<,...,<y_{k}$ such that for each training instance $x_{l}$ it can be associated a label $y_{l}.$ The problem is to determine a ranking function that orders a new set of instances $\{x^{\prime}_{j}\}^{t}_{j=1}$ according to their (unknown) preference degrees. The performance meaures for this are: AUC($k=2$, for the bipartite ranking) and C-index in the polytomous case ($k>2$, for $k$-partite ranking). As now the ranker has the ability to reject predictions, there is a trade-off between correctness and completeness. Correctness is measured by gamma rank correlation[269] is a measure of rank correlation, i.e., the similarity of the orderings of the data when ranked by each of the quantities. Completeness measure punishes the abstention from comparisons that should actually be made. Furthermore, a preference relation[268] $P:A\times A\rightarrow\left[0,1\right]$ provides a measure of support for the pairwise preference $a\succ b$ with $P(a,b)=1-P(b,a)$ for all $a,b\in A$ which can be considered an application of a generic approach to transform every ranker into a partial ranker via ensembling such that
@@
    For any ranker $L$, train $k$ ranking models $M_{1}…M_{k}$ by resampling from the original data set, i.e., by $k$ bootstrap samples. By querying these models, $k$ rankings $\succ_{1}...\succ_{k}$ are generated.
@@ 3.9 Machine learning using formal concept analysis
Formal concept analysis(FCA) [7] is a well known method which stems directly from partial order and lattice theory for conceptual knowledge representation and data analysis in information science. Formal concept analysis was introduced in the seminal work of Wille[181] in 1982 as a result of an attempt to restructure mathematical order and lattice theory. The fundamental theory underlying FCA is the representability of complete lattices by ordered sets of their meet- and join-irreducibles [231]. Since ordered sets of irreducibles are naturally represented by binary matrices, this makes it possible to apply certain aspects of the lattice theory to the analysis of data given by object-attribute matrices.[198]. Furthermore, the row of the binary matrix is usually interpreted as an object and a column is interpreted as a binary attribute. If an object possesses an attribute the corresponding entry is 1 and if it does not possess it then the correesponding entry is 0 [298]. It has also been shown that through a process of discretising and booleanising data, it is possible to convert a variety of data sets into formal contexts or concept lattices [278, 279, 280]. FCA can be viewed as an unsupervised machine learning technique. In this direction, it takes an input binary relation which is represented by a binary matrix. Its goal is usually to determine the natural concepts described in the data, and then organizes the concepts in the form of a Hasse diagram corresponding to a partial ordering [70, 316, 317]. Due to its generality, simplicity, and powerful mathematical foundation, there is currently a growing adoption of FCA in a wide variety of data science tasks. Its advantage can be seen in the several FCA based machine learning methods that gave competitive results in various machine learning problems compared to classical methods [185]. In a recent study on the application of formal concept analysis in image classification problems, Khatri[184] establishes several advantages of formal concept analysis over convolutional neural networks for image classification task. These advantages are :
@@
  • Firstly, FCA provides clarity which can be useful for interpretability of machine learning classification tasks. Furthermore, FCA facilitates examination of classification hierarchy by providing a visual representation of data in the form of a lattice.
@@
  • Secondly, during the classification process, data can be added and removed from the lattice without the need for retraining. Moreso, the FCA classifier outperforms the Random forest classifier, which is a human interpretable model.
An important feature of FCA-based classification methods is that they do not make any assumptions regarding statistical models of a dataset[216]. A typical refinement of the idea of learning is learning from positive and negative examples, when a learning system obtains descriptions of positive and negative examples and constructs a generalization of positive examples that does not “cover” negative examples[198]. There are several FCA-based models for classification based on learning from positive and negative examples [187], [197],[198]. In [233] a genetic algorithm based machine learning method is presented for automatically learning object-oriented hierarchies very similar to lattice based concept hierarchies from large data set. Ikeda and Yamamoto[239] propose a method for classifying data that generates a concept lattice and selects appropriate formal concepts in the lattice. The method is useful for robust feature selection during classification by securing sufficient storage space in order to maintain both the selected concepts and redundant concepts. This is beneficial in terms of reduced time duration in solving practical problems related to a variety of large data sets. Formal concept analysis has been applied in the context of ensemble learning: Dagging [267, 256], Recommendatory-based Multiple Classifier System [266]
@@
FCA based approaches have also been used for the automated discovery of domain-specific ontologies from textual descriptions of domain entities [318, 319]. Ontology learning[320] refers to the process of automatically extracting and constructing knowledge structures or models from unstructured or semi-structured data sources such as text, speech, images, or sensor measurements. These knowledge structures typically take the form of annotated taxonomies, concept hierarchies, or domain-specific ontologies that capture various aspects of the underlying domain or subject matter. Ontologies are a key element of the semantic web. They aim to capture basic knowledge by providing appropriate terms and formal relationships between them, so that they can be used in a machine-processable manner. Relational Concept Analysis (RCA) is an extension of the FCA framework to take into account multi-relational datasets, RCA generates a family of concept lattices, precisely one for each category of objects Using formal concept analysis as well as relational concept analysis different approaches have been implemented for ontology learning and ontology ex-traction [234, 235, 236, 237, 232]. Many machine learning based AI systems have been designed as black boxes. which makes it challenging in achieving interpretability in AI models due to the complexity of AI algorithms. Sangroya et al.[281] propose a general concept lattice theory based framework for explainability of AI models. The main idea is that given an outcome of a deep learning model and a domain ontology, the objective of their proposed solution is to identify an explanation that can point the user to the prominent feature set for a certain outcome.

### v2

@@ 3.3 Natural Language Processing
In natural language processing, compositionality[148] describes the ability to easily recombine familiar units like words into new phrases and sentences. In recent years, this has been the focus of intense interest in artificial intelligence. However, existing neural network based models have proven to be extremely deficient in such a capability[105]. In order to improve on the compositional generalization ability of neural encoder-decoder architectures for natural language understanding, Guo et al.[150] introduce a hierarchical poset decoding paradigm with the goal of taking into consideration the partial permutation invariance and the poset structure of semantics. Compositional generalization is important in tasks that can involve complex natural language utterances, which can be associated to many equivalent meaning representations. One of such task is dependency parsing. Semantic dependency parsing[199] is the task of mapping natural language sentences into representations of their meaning in the form of directed graphs on words. Dependency parsing is generally used to identify semantic relations between words in a sentence by analyzing their grammatical structure. Dyer [116] proposes a method for converting dependency trees to surface orders using syntactic word embeddings and edge-weighted posets. Graph neural network were employed to learn the edge weights of the poset representing the surface order, and the training was carried out on the Universal Dependencies (UD) corpora.
@@ 3.4 Classification
Most major studies in artificial intelligence has significantly concentrated on developing methods for classification. Classification in machine learning involves prediction tasks on discrete target variable. It is common to find learning tasks that involve ranking of objects, particularly in domains that generate relational data where the goal of inferring is conditioned on a particular target object. In classification, a single label $y\in Y$ can be associated with a covariate $x$. A generalization of this problem is conditional ranking, the task of assigning to $x$ a full or partial ranking of the items in Y. Lebanon and Lafferty[204] propose a unifying algebraic framework for ensemble methods of classification and conditional ranking, focusing on the case of discrete and ranked inputs. Their approach is based on the ranking poset having a fixed size of items, and which consists of the collection of all full and partial rankings endowed with a partial order given by refinement of rankings. The structure of the poset of partial ranking over Y gives rise to natural invariant distance functions which generalizes Kendall’s Tau[188] and the Hamming distance. Using the invariance properties of the distances, a generative interpretation of the probabilistic model is derived, and this is suggested to be useful in model selection and validation[204]. Zhang[254] presents a framework for ranking with classification rules. Classification rules can be generated using inductive learning algorithms[6] on data. Sahami[254] introduces an inductive machine learning algorithm called Ruleamer, which generates a set of symbolic classification rules as output using lattices. The Ruleamer algorithm takes as input (i) a lattice, L, (ii) a set of instance classification labelings, C, which corresponds to the instance nodes in L, and (iii) a noise parameter, N, indicating a percentage by which each induced rule can misclassify some portion of the training instances to which it applies. Learning classification rules has also been extended to sequence data [102, 101]. The problem of sequential classification is considered when classification states form lattices and response distributions are experiment-specific. Tatsuoka[278] proposes a Bayesian framework for sequential classification on finite lattice models in which response distributions are allowed to vary according to experiment. When the models are latent and complex, such as in cognitive applications, it becomes imperative to have available a variety of data analytic tools for fitting the models, and for the validation of assumptions that are made regarding the specification of class conditional response distributions for experiments. A framework for model fitting and analysis of experiments has been developed byTatsuoka[272]. This led to the formulation of the theory on asymptotically optimal sequential selection of experiments in [273, 274], which can be applied to Bayesian classification problems when the parameter space is finitely partially ordered. In general, lattices can be used as classification models when it is natural to consider classification states as partially ordered. Examples include cognitive modeling, such as in neuropsychological (NP) assessment, educational testing, and in group testing data.
An important task involving supervised machine learning techniques is classifying content employed in many contexts. Multi-label classification[36] deals with the problem in which each instance can be associated with a set of class labels. But a drawback on the application of multilabel classification in many real-world applications arises when classes overlap in the feature space or the number of class labels becomes very large resulting to an exponentially growing number of label combinations. As such it becomes computationally inefficient to use most multi-label classification algorithms in such scenarios. Multilabel classification has been suggested as a powerful tool for building AI systems that can handle complex uncertain scenarios. Uncertainty is a pervasive aspect of many AI systems, as they often deal with incomplete or conflicting information caused by the variability in data and models. The mathematical theory of belief functions as applied in the Dempster-Shafer theory(DST) introduced in [260, 108], provides a generalized framework which blends the concept of probability associated with uncertainty and the notion of sets related to logic. However, the computational complexity of DST increases significantly when confronted with a substantial number of events or sources of evidence, resulting in potential performance challenges caused by the exponential complexity of it’s operations. In addition, problems which involve learning instances belonging to several classes(frame of discernment) simultaneously which may pose several difficulties in obtaining the right label combination during classification. Denoeux[109] suggests that when the frame of discernment has a lattice structure, it is possible to restrict the set of events to intervals in that lattice. This approach involving lattices leads to the application of Dempster-Shafer theory to computationally demanding machine Learning tasks such as multi-label classification and ensemble clustering.
3.5 Deep Learning
@@
In this section, the use of posets in deep unsupervised learning, attention-based neural networks, and topological deep learning are reviewed. Deep unsupervised learning[186] is an emerging discipline in the field of artificial intelligence and machine learning. It can serve as an alternative to supervised approaches in representation learning. In particular, unlabelled data is much cheaper to obtain which is a key advantage when adopting deep unsupervised learning for complex processing tasks involving large datasets. One of such task is visual similarities learning[95] in the field of computer vision. Various approaches have been adopted for visual similarity learning such as the use of convolution neural networks(CNN) [294]. But this approach has a major limitation in that it requires millions of samples of supervised training in order to achieve a good performance. Bautista et al.[31], propose an unsupervised approach to visual similarity learning based on CNNs by framing it as a combination of surrogatee (i.e. artificially created) classification tasks and poset ordering. This approach which combines classification with a partial ordering of samples resolved some of the major problems of visual similarity learning such as:
@@
Neural attention models[222] focus on specific regions of images instead of the entire picture, originated in an attempt to reduce the computational complexity of image processing while improving performance. In general, the application of attention neural models in the field of computer vision was motivated from the idea of mimicking human attention. Many different neural attention models[256, 41] are now available and have been a very active research area over the past decade. Attention is an important mechanism that can be employed for a variety of deep learning models across many different domains and tasks. In natural language processing NLP, transformers and attention have been utilized successfully in a plethora of tasks including reading comprehension, abstractive summarization, word completion, and so forth. Hajij et al.[154] introduces higher-order attention networks (HOANs), as a special class of attention-based neural networks defined on a generalized higher-order domain called a combinatorial complex (CC). CCs define a structure that bridges the gap between simplicial/cell complexes and hypergraphs. In addition, CC is a poset; i.e., a partially ordered set with partial order relation given by set inclusion. Specifically, Hajij et al.[154] demonstrate the reducibility of any CC to a Hasse graph, allowing the characterization of certain computational and conceptual aspects of HOANs in terms of graph-based models. Furthermore, the Hasse graph is used to provide a definition of equivarence for HOANs. The developments of HOANs subsequently influenced the study of topological deep learning with respect to combinatorial complexes. Topological Deep Learning (TDL)[311] is a rapidly evolving field which provides a comprehensive framework that combines the principles of topological data analysis (TDA) with deep learning techniques. In particular, it uses topological features to understand and design deep learning models. TDL originated from the pioneering works on Topological Signal Processing (TSP)[26] that showed the benefits of considering higher-order (multi-way) relationships among data points. The growing interest driving the current developments in topological deep learning is attributed to the fact that many systems are characterized by higher-order interactions that cannot be captured by the intrinsically pairwise structure of graphs. Recently, Hajij et al. [155] developed a general TDL framework along with the introduction of an abstract class of neural networks called combinatorial complex neural networks (CCNNs). CCNNs can also be viewed as a framework that generalizes many popular architectures, such as convolutional and attention-based neural networks. A CC-based neural network is reducible to graph-based models and is capable of exploiting all neighborhood matrices or a subset of them. The CC structure class is determined by the incidence matrices, adjacency matrices, coadjacency matrices(Proposition 8.1.[155]). Combinatorial complexes (CCs)[30] are able to handle both hierarchical and set-type higher-order interactions. Hasse graph of the CC, describes the poset structure between cells. The structure of combinatorial complexes permits graph-based pooling, facilitate message passing of topological features in deep learning, flexible modeling of relations among relations. Based on this structural benefits, Hajij et al.[155], present the Hasse graph interpretation of CCNNs as follows. (i) It is shown that any CCNN-based computational model can be realized as a message-passing scheme over a subgraph of the augmented Hasse graph of the underlying CC; (ii) A tensor diagram represents a CCNN via a directed graph. Any computation on tensor diagrams CCNN is realizable on augmented Hasse graphs.
@@ 3.6 Semi-Supervised Learning
Semi-supervised learning[286] is a branch of machine learning that bridges supervised and unsupervised learning by combining both labelled and unlabelled data to train artificial intelligence (AI) models for classification or regression. Semi-supervised classification methods are particularly relevant in scenarios where labelled data is scarce, expensive or difficult to obtain. Graph-based semi-supervised learning methods[264] have demonstrated their benefits in various domains due to their uniqueness of structure, the universality of applications, and their scalability to large scale data. Graph-based learning methods explicitly considers the relations between two entities (i.e.vertices). In an ordinary graph an edge connects exactly two vertices. On the other hand, a hypergraph represents a generalization of a graph in which an edge can join any number of vertices. But in most settings, hypergraph structure represents the hyperedge as an unordered set of vertices without taking into consideration the possibility of ordering relationships among vertices which is inherent in most real-world data. This motivated Feng et al. [132] to suggest two related solutions which are:
  • the construction of a new data structure named Partial-Order Hypergraph, which specifically injects the partial ordering relations among vertices into a hyperedge ;
@@
Furthermore, graph convolution network(GCN)[191] was applied in a semi-supervied setting for learning over the constructed partial-order hypergraphs.
@@ 3.7 Time Series Modeling
Many machine learning and statistical models have been suggested for time series modeling of sequential data [302]. One way of generating sequence data is from a series of discrete events that occur over time. Event sequence data arise in many applications, such as web browsing, e-commerce, process monitoring etc. An important problem in mining sets of sequences of events is to get an overview of the ordering relationships in the data. Mannila & Meek[212] introduce a method for discovering partial orders from a sequence of events. The main idea of the method is to view a partial order as a generative model for sequences and by describing a set of sequences by means of mixture models of partial orders. In addition, it is also shown that the likelihood of a given partial order is inversely proportional to the number of total orders compatible with the partial order. The computation of the number of linear extensions is restricted to series parallel posets which are computationally efficient. A learning algorithm is constructed based on greedy search over partial orders defined on the set of possible events. This learning algorithm determines if a sequence is compatible with a partial order in linear time. Nicholls et al [228] explore the Bayesian inference for partial orders from random linear extensions using time-series data of social hierarchy covering the time period between the $11^{\text{th}}$ and $12^{\text{th}}$ centuries. This approach introduces a new class of models for rank-order time series data in which actors are listed in order of precedence which is presented as a collection of lists. The lists are modeled as a realisation of a queue in which queue-position is constrained by an underlying social hierarchy corresponding to a partial order, which is used in modelling the evolution of the social hierachy over the given time period..
@@ 3.8 Learning to rank
Learning to rank (LTR)[70, 308] describes a class of algorithmic techniques that apply supervised machine learning to solve ranking problems from a wide range of domains such as information retrieval, recommendation systems, search engine optimization etc. It arises from the need to obtain prediction results based on the best order of their relevance to a machine learning classification problem. Several methods have been proposed for LTR and they can be categorized or grouped into three main approaches: pointwise, pairwise , listwise. Pointwise approach involves scoring items independently and then ranking them based on their scores. In contrast the pairwise or listwise ranking methods, consider the relative positions of items in pairs or lists respectively. The goal for the ranker is to minimize the number of inversions in ranking i.e. cases in which the pair of results are in the wrong order relative to the ground truth. In particular, the listwise learning approach addresses the ranking problem, by analyzing ranked lists of objects as input instances and then trains a ranking function through the minimization of a listwise loss function defined on the predicted list and the ground truth list. In [96], it is noted that not all rankings can follow a strict total order in the context of learning to rank. As such a proposed method of relaxing the conventional setting such that predictions are given in terms of partial instead of total orders is outlined. The key idea is that if a model is uncertain about the relative order of two alternatives, in which case it is unable to clearly determine whether the former should precede the latter and vice-versa, it may decline or pospone the making of a conclusive decision and instead declare such pair of alternatives as incomparable. In general, learning to rank problems as presented in [97] can be formulated as follows : Given a set of training instances $\{x_{1},...,x_{n}\}\subseteq\mathcal{X}$ and a set of labels $\mathcal{Y}=\{y_{1},...,y_{k}\}$ endowed with an order $y_{1}<y_{2}<,...,<y_{k}$ such that for each training instance $x_{l}$ it can be associated a label $y_{l}.$ The problem is to determine a ranking function that orders a new set of instances $\{x^{\prime}_{j}\}^{t}_{j=1}$ according to their (unknown) preference degrees. The performance meaures for this are: AUC($k=2$, for the bipartite ranking) and C-index in the polytomous case ($k>2$, for $k$-partite ranking). Since the ranker has the ability to reject predictions, there is a trade-off between correctness and completeness. Correctness is measured by gamma rank correlation[249] is a measure of rank correlation, i.e., the similarity of the orderings of the data when ranked by each of the quantities. Completeness measure penalizes the abstention from comparisons that should actually be made. Furthermore, a preference relation $P:A\times A\rightarrow\left[0,1\right]$ provides a measure of support for the pairwise preference $a\succ b$ with $P(a,b)=1-P(b,a)$ for all $a,b\in A,$ which can be considered an application of a generic approach[97] to transform every ranker into a partial ranker via ensembling such that:
@@
    for any ranker $L$, train $k$ ranking models $M_{1}…M_{k}$ by resampling from the original data set, i.e., by $k$ bootstrap samples. By querying these models, $k$ rankings $\succ_{1}...\succ_{k}$ are generated;
@@ 3.9 Machine learning using formal concept analysis
Formal concept analysis(FCA) [11, 141] is a well known method which originates from partial order and lattice theory for the purpose of conceptual knowledge representation and data analysis in information science. Formal concept analysis was introduced in the seminal work of Wille(1982) as a result of an attempt to restructure mathematical order and lattice theory. The basic data format in FCA[248] is a cross-table given by the triple $\left(O,A,I\right)$ called formal context , where $O$ is a set of elements(referred to as formal objects), $A$ a set of properties (called formal attributes) and $I$ the relation has on $O\times A.$ A pair $\left(X,Y\right)$ where $X$ is a maximal set of elements(called extent) and $Y$ is a maximal set of shared properties (called intent) represents a formal concept. The set of all concepts of the context is partially ordered by extent inclusion which represents a complete lattice. The fundamental theory underlying FCA is the representability of complete lattices by ordered sets of their meet- and join-irreducibles [27]. Since ordered sets of irreducibles are naturally represented by binary matrices, this makes it possible to apply some properties of lattice theory to the analysis of data given by object-attribute matrices[201]. Furthermore, the rows of the binary matrix are usually interpreted as an object and a column is interpreted as a binary attribute. If an object possesses an attribute the corresponding entry is 1 and if it does not possess it then the corresponding entry is 0 [200]. It has also been shown that through a process of discretising and booleanising data, it is possible to convert a variety of datasets into formal contexts or concept lattices [9, 8, 10]. FCA can also be viewed as an unsupervised machine learning technique. In this direction, it takes an input binary relation which is represented by a binary matrix. Its goal is usually to determine the natural concepts described in the data, and then organizes the concepts in the form of a Hasse diagram corresponding to a partial ordering [91, 143, 164]. Due to its generality, simplicity, and powerful mathematical foundation, there is currently a growing adoption of FCA in a wide variety of data science tasks[202]. Its advantage can be seen in the several FCA based machine learning methods that gave competitive results in various machine learning problems compared to classical methods [5]. In a recent study on the application of formal concept analysis in image classification problems, Khatri[196] establishes several advantages of formal concept analysis over convolutional neural networks for image classification task. These advantages are :
@@
  • FCA provides clarity which can be useful for interpretability of machine learning classification tasks. Furthermore, FCA facilitates examination of classification hierarchy by providing a visual representation of data in the form of a lattice.
@@
  • During the classification process, data can be added and removed from the lattice without the need for retraining. Moreso, the FCA classifier outperforms the random forest classifier, which is a human interpretable model.
@@
An important feature of FCA-based classification methods is that they do not make any assumptions regarding statistical models of a dataset[237]. A typical refinement of the idea of learning, is learning from positive and negative examples, when a learning system obtains descriptions of positive and negative examples and constructs a generalization of positive examples that does not “cover” negative examples[201]. There are several FCA-based models for classification based on learning from positive and negative examples [202], [229],[201]. In Jabin[170] a genetic algorithm based machine learning method is presented for automatically learning object-oriented hierarchies which are very similar to lattice based concept hierarchies from large datasets. Ikeda and Yamamoto[165] propose a method for classifying data that generates a concept lattice and selects appropriate formal concepts in the lattice. The method is useful for robust feature selection during classification by securing sufficient storage space in order to maintain both the selected concepts and redundant concepts. This is beneficial in terms of reduced time duration in solving practical problems related to a variety of large datasets. Formal concept analysis has been applied in the context of ensemble learning: Dagging [218, 5], Boosting[217], Bagging [184], Recommendatory-based Multiple Classifier System [187]. In Xie[301], concept lattice(CL) model, the core structure in formal concept analysis integrates the simple base classifer (Naive Bayes(NB) and Nearest Neighbour(NN)) into each node of the concept lattice to form a new composite classifer namely- Concept Lattice Naive Bayes(CLNB) and Concept Lattice Nearest Neighbour(CLNN) classification systems. Experimental results on 26 datasets indicate that two hybrid classifcation systems(CLNN and CLNB) perform better than their corresponding base classifers and CLNB even outperforms state-of-the-art classifiers. Many machine learning based AI systems have been designed as black boxes. which makes it challenging in achieving interpretability in AI models due to the complexity of AI algorithms. Sangroya et al.[255] propose a general concept lattice theory based framework for explainability of AI models. The main idea is that given an outcome of a deep learning model and a domain ontology, the objective of their proposed solution is to identify an explanation that can point the user to the prominent feature set for a certain outcome. FCA based approaches have also been used for the automated discovery of domain-specific ontologies from textual descriptions of domain entities [25, 283, 310]. Ontology learning[225] refers to the process of automatically extracting and constructing knowledge structures or models from unstructured or semi-structured data sources such as text, speech, images, or sensor measurements etc. These knowledge structures typically take the form of annotated taxonomies, concept hierarchies, or domain-specific ontologies that capture various aspects of the underlying domain or subject matter. Ontologies are a key element of the semantic web. They aim to capture basic knowledge by providing appropriate terms and formal relationships between them, so that they can be used in a machine-processable manner. Relational Concept Analysis (RCA)[163] is an extension of the FCA framework to take into account multi-relational datasets, RCA generates a family of concept lattices, precisely one for each category of objects Using formal concept analysis as well as relational concept analysis different approaches[159, 179, 153, 230, 100] have been implemented for ontology learning and ontology extraction.